OpenStudy (anonymous):

Is sinx cosx ever equal to 1?

OpenStudy (unklerhaukus):

\[\sin(x)\times\cos(x)≠1\]

OpenStudy (unklerhaukus):

to get a product equal to one , either both must be 1 or one is less than 1 and the other greater, sine and cosine are always ≤1

OpenStudy (experimentx):

let us suppose sin(x)cos(x) = 1 then, 2 sin(x) cos(x) = 2 sin(2x) = 2 since ... sine of any value is not greater than 1, our assumption must be wrong